# First-Mover Bias in Gradient Boosting Explanations: Mechanism, Detection, and Resolution

### 22 Mar 2026 v1

@@ 6.2 The principle: independence largely resolves it
0.7 DASH vs SB Equity 0.018 -1.07
0.7 DASH vs LSM Equity <<0.001 -1.96
@@
0.9 DASH vs SB Equity 0.015 -1.27
0.9 DASH vs LSM Equity <<0.001 -3.03
@@
The SR equivalence is the headline theoretical result.
We find that model independence alone—even in its simplest form (seed averaging, no diversity optimization)—largely resolves first-mover bias in the linear regime. Stochastic Retrain achieves marginally higher stability point estimates than DASH at most $\rho$ levels (0.9750.9750.9800.980 vs. 0.9720.9720.9770.977), and the non-significant differences on accuracy ($d=+0.05$, $p=0.926$) and equity (d=0.21d=-0.21, $p=0.622$) confirm this equivalence statistically. This equivalence is our central theoretical contribution: it isolates model independence as the operative mechanism, and positions DASH as a principled instantiation of that principle rather than a statistically superior pipeline. Phrased differently: raw seed averaging already delivers most of the benefit; DASH then operationalizes this principle with engineered quality control (filtering, diversity-aware selection), diagnostic tools, and hyperparameter robustness that seed averaging cannot provide.
@@
    Equity. Marginally lower within-group CV ($0.176$ vs. $0.182$), suggesting that forced feature restriction distributes credit more evenly, though this difference is not statistically significant and should be interpreted cautiously.
@@ 6.3 The effect scales with correlation
Table 3 (Section 6.1) reveals a dose-response relationship: as $\rho$ increases, first-mover bias intensifies for dependent methods while independent methods remain immune. LSM stability degrades monotonically from $0.953$ at $\rho=0$ to $0.925$ at $\rho=0.95$—a $2.9\%$ decline. Single Best follows the same pattern ($0.973\to 0.951$, a $2.3\%$ decline). DASH’s stability is effectively flat (0.9720.9720.9770.977), fluctuating by less than $0.5\%$ across the entire correlation range.
The equity metric shows the same scaling. LSM’s within-group CV worsens from $0.170$ to $0.284$ ($67\%$), while DASH ranges from $0.164$ to $0.176$ ($7\%$). This is consistent with sequential residual dependency concentrating importance within correlated groups, with the concentration scaling with the degree of correlation—as the first-mover bias hypothesis predicts. When correlation is low, features within a group carry sufficiently distinct signal that the first-mover advantage is weak. When correlation is high, the features are near-interchangeable, and the first split’s arbitrary selection dominates.
At $\rho=0$, Single Best, Stochastic Retrain, and DASH cluster tightly (stability 0.9720.9720.9750.975), satisfying the safety desideratum: DASH does not degrade explanations when multicollinearity is absent. LSM already trails at $0.953\pm 0.003$, and its gap widens monotonically with $\rho$.
@@
Figure 3: Stability, accuracy, and equity as a function of within-group correlation $\rho$ (linear DGP, 20 repetitions per level). Independent methods (DASH, Stochastic Retrain) are flat across correlation levels; dependent methods (Single Best, LSM) degrade monotonically. All seven methods shown; Table 3 reports the four principal methods in detail, Table 4 compares all at $\rho=0.9$.
@@ 6.4 Detecting first-mover bias: FSI and IS Plot
A key practical question is: how can a practitioner know whether their explanations suffer from first-mover bias? Ground-truth importance is never available in practice. DASH’s Stage 5 diagnostics address this by quantifying explanation disagreement across the ensemble.
@@ 6.5 Real-world validation
On the Breast Cancer dataset (30 features, 21 pairs with $|r|>0.9$, 20 repetitions), DASH nearly triples stability:
@@
Breast Cancer Single Best (M=200M{=}200) $317\pm.053$
DASH (MaxMin) $\mathbf{.930}\pm.005$ +0.613+0.613
Superconductor Single Best $830\pm.020$ $9.18\pm 0.11$
Large Single Model $689\pm.030$ 0.141-0.141 $9.34\pm 0.09$
DASH (MaxMin) $\mathbf{.962}\pm.010$ +0.132+0.132 $\mathbf{9.15}\pm 0.08$
Calif. Housing Single Best $967\pm.010$ $0.460\pm 0.009$
DASH (MaxMin) $\mathbf{.982}\pm.005$ +0.015+0.015 $\mathbf{0.450}\pm 0.005$
The Breast Cancer improvement is the largest across all experiments. The training-budget-matched Single Best (M=200M{=}200) achieves only $0.317$ stability, because Breast Cancer’s extreme collinearity (21 pairs with $|r|>0.9$) makes model selection itself unstable—with 200 hyperparameter configurations, the “best” model varies wildly across repetitions, and feature rankings are highly unstable across runs. The radius/perimeter/area triad creates extreme SHAP instability: a feature’s importance can swing from first to fifth across runs. DASH consensus substantially stabilizes this, producing stable top features aligned with known geometric redundancy in the feature set—mean concave points (0.228), worst area (0.201), worst perimeter (0.201)—that are stable across repetitions.
@@
On the Superconductor dataset (81 features, 21,263 samples), DASH improves stability by +0.132+0.132 over Single Best and +0.273+0.273 over the Large Single Model, while achieving marginally better predictive RMSE. The LSM result is particularly striking: despite having the same total tree count as DASH, its sequential training produces the poorest reproducibility—consistent with first-mover bias at scale.
@@
On California Housing (8 features, 20,640 samples), several feature pairs have $|r|>0.7$ (e.g., latitude/longitude, income/house value). DASH improves stability by +0.015+0.015 over Single Best while maintaining equivalent predictive RMSE. The improvement is smaller than for Breast Cancer and Superconductor, consistent with the milder degree of collinearity and the small feature space.
@@ 6.6 Robustness and scope
Under the nonlinear DGP (Table 7), all methods degrade (stability drops from $\sim$0.93 to $\sim$0.88), but the first-mover bias mechanism still operates: DASH outperforms Single Best at $\rho\geq 0.7$ (stability gap +0.078+0.078 at $\rho=0.95$).
Table 7: Nonlinear DGP: stability and equity across correlation levels. SB = Single Best, LSM = Large Single Model, LSM-T = LSM (Tuned), SR = Stochastic Retrain. Five methods with distinct training structures are shown; Random Selection and Naive Top-$N$ (which performed similarly to DASH in the linear regime) are omitted for space and are available in the code repository. Bold indicates best per $\rho$.
@@
0.0 0.178 0.181 0.181 0.180 0.168
0.5 0.177 0.189 0.174 0.169 0.172
0.7 0.192 0.203 0.184 0.180 0.170
0.9 0.200 0.241 0.220 0.177 0.173
0.95 0.231 0.267 0.229 0.176 0.167
At $\rho=0$, DASH and Single Best perform nearly identically ($0.934$ vs. $0.933$). At $\rho=0.5$, Stochastic Retrain achieves the highest stability ($0.855$), with DASH ($0.852$) marginally ahead of Single Best ($0.849$). DASH’s advantage over Single Best grows consistently with correlation. This is a genuine scope boundary: nonlinear relationships are common in practice, and overall stability levels are lower than for the linear DGP ($\sim$0.87–0.93 vs. $\sim$0.93–0.98). When the DGP contains interactions and nonlinear terms, different models in the DASH ensemble may capture different interaction structures. Averaging SHAP values across models that have learned qualitatively different functional forms introduces noise rather than canceling arbitrary choices.
At $\rho\geq 0.7$, first-mover bias reasserts itself as the dominant source of instability, and DASH’s independence-based cancellation provides clear benefit (+0.078+0.078 at $\rho=0.95$). Practitioners should therefore check both the degree of collinearity and the presence of strong nonlinear interactions before applying DASH. The FSI diagnostic (Section 6.4) can help: if FSI values are low across all features, ensemble averaging may not be needed.
@@ 7.1 Reframing explanation instability
This work reframes the SHAP instability problem from “SHAP is noisy under multicollinearity” to a specific mechanistic hypothesis: sequential residual dependency in gradient boosting creates a first-mover bias that concentrates feature importance on arbitrary features within correlated groups. Three lines of evidence support this view: (1) the Large Single Model—which maximizes sequential dependency—produces the poorest reproducibility at every correlation level; (2) all methods that achieve model independence restore stability to the same level ($\approx 0.977$ at $\rho=0.9$), regardless of their other design choices; and (3) the effect scales with correlation severity, as the hypothesis predicts.
@@
The DASH ensemble’s $K$ importance vectors also support a richer output representation: features that consistently swap rank across ensemble members can be treated as incomparable rather than arbitrarily ordered, yielding a partial order on feature importance that more honestly reflects genuine attribution ambiguity under collinearity. We pursue this direction in companion work.

### 06 Apr 2026 v2

@@ 6.2 The principle: independence largely resolves it
0.7 DASH vs SB Equity <<0.001 -1.03
0.7 DASH vs LSM Equity <<0.001 -1.78
@@
0.9 DASH vs SB Equity <<0.001 -1.47
0.9 DASH vs LSM Equity <<0.001 -2.94
@@
The SR equivalence isolates the operative mechanism.
Stochastic Retrain implements the same core operation as DASH—averaging attributions across independently trained models—differing only in how independence is achieved (seed diversity vs. hyperparameter diversity). The companion paper (Caraker et al., 2026b) proves this averaging is the unique minimum-variance unbiased estimator; the equivalence below confirms that practical independence, not pipeline engineering, drives stability. We find that model independence alone—even in its simplest form (seed averaging, no diversity optimization)—largely resolves first-mover bias in the linear regime. Stochastic Retrain achieves marginally higher stability point estimates than DASH at most $\rho$ levels (0.9750.9750.9800.980 vs. 0.9730.9730.9770.977), and the non-significant differences on accuracy ($d=+0.05$, $p=0.926$) and equity (d=0.13d=-0.13, $p=0.44$) confirm this equivalence statistically. This equivalence is the strongest evidence for our central claim: that model independence, not any particular aggregation strategy, is the operative mechanism that neutralizes first-mover bias. Phrased differently: raw seed averaging already delivers stability in the linear regime; DASH then operationalizes this principle with additional benefits—nonlinear robustness (Section 6.6), equitable attribution (below), ground-truth-free diagnostics (Section 6.4), and computational efficiency—that simple seed averaging does not provide. The companion paper (Caraker et al., 2026b) proves that DASH consensus averaging is Pareto-optimal among all attribution aggregation methods: it achieves the Cramér–Rao variance bound $\sigma^{2}/M$ and no method can simultaneously achieve zero within-group unfaithfulness and higher between-group stability for the same ensemble size.
@@
    Equity. Significantly lower within-group CV than Random Selection ($0.176$ vs. $0.187$; $p<0.001$, d=0.37d=-0.37), indicating that forced feature restriction distributes credit more evenly across correlated groups. The equity advantage over SR is smaller and not significant (d=0.13d=-0.13, $p=0.44$).
Beyond stability: nonlinear robustness.
The linear regime demonstrates that any form of model independence suffices for stability. Section 6.6 shows that under nonlinear data-generating processes, how independence is achieved matters: DASH’s population-level diversity—forced feature restriction and varied hyperparameters—outperforms seed averaging by +0.030+0.030 stability at $\rho=0.9$ (Table 9), likely because diverse feature subsets explore distinct nonlinear interaction pathways that seed averaging, which fixes hyperparameters, cannot reach.
Top-$k$ ranking stability.
Beyond overall rank correlation, we examine the stability of the top-$5$ feature subset across repetitions (top-$k$ overlap). SR achieves higher top-$5$ ranking stability than DASH ($0.922$ vs. $0.863$ at $\rho=0.9$; $p=0.18$, n.s.), likely because fixed hyperparameters preserve the feature importance landscape across seeds while DASH’s diversity mechanism produces more varied individual-model rankings. However, DASH significantly outperforms all dependent methods on top-$k5$: +0.317+0.317 vs. Single Best, +0.430+0.430 vs. LSM, and +0.488+0.488 vs. Random Forest (all $p<0.001$). Diversity selection also helps: DASH exceeds Random Selection by +0.036+0.036 on top-$k5$, indicating that MaxMin selection improves not only equity but also the consistency of top-feature identification.
@@ 6.3 The effect scales with correlation
Table 3 (Section 6.1) reveals a dose-response relationship: as $\rho$ increases, first-mover bias intensifies for dependent methods while independent methods remain immune. LSM stability degrades monotonically from $0.955$ at $\rho=0$ to $0.927$ at $\rho=0.95$—a $2.9\%$ decline. Single Best follows the same pattern ($0.972\to 0.952$, a $2.1\%$ decline). DASH’s stability is effectively flat (0.9730.9730.9770.977), fluctuating by less than $0.5\%$ across the entire correlation range.
The equity metric shows the same scaling. LSM’s within-group CV worsens from $0.170$ to $0.277$ ($67\%$), while DASH ranges from $0.164$ to $0.176$ ($7\%$). This is consistent with sequential residual dependency concentrating importance within correlated groups, with the concentration scaling with the degree of correlation—as the first-mover bias hypothesis predicts. When correlation is low, features within a group carry sufficiently distinct signal that the first-mover advantage is weak. When correlation is high, the features are near-interchangeable, and the first split’s arbitrary selection dominates.
At $\rho=0$, all independent methods achieve stability $\geq 0.972$, with differences $\leq 0.003$. Stochastic Retrain is marginally higher ($0.975$ vs. DASH’s $0.973$), a statistically detectable difference ($p<0.001$ by bootstrap test) but practically negligible—two orders of magnitude below the $0.1$ safety threshold. This satisfies the safety desideratum: DASH does not meaningfully degrade explanations when multicollinearity is absent. Notably, the SR–DASH gap narrows from $0.002$ at $\rho=0$ (significant) to <0.001<0.001 at $\rho=0.9$ (n.s.), consistent with diversity selection becoming more valuable as correlation intensifies. LSM already trails at $0.955\pm 0.003$, and its gap widens monotonically with $\rho$.
@@
Figure 4: Stability, accuracy, and equity as a function of within-group correlation $\rho$ (linear DGP, 50 repetitions per level). Independent methods (DASH, Stochastic Retrain) are flat across correlation levels; dependent methods (Single Best, LSM) degrade monotonically. All seven methods shown; Table 3 reports the four principal methods in detail, Table 4 compares all at $\rho=0.9$.
@@ 6.4 Detecting first-mover bias: FSI and IS Plot
A key practical question is: how can a practitioner know whether their explanations suffer from first-mover bias? Ground-truth importance is never available in practice. DASH’s Stage 5 diagnostics address this by quantifying explanation disagreement across the ensemble. The companion paper (Caraker et al., 2026b) introduces complementary formal tests: a single-model split-frequency $Z$-test (F5: tests whether the split-count ratio within a collinear pair deviates from the null of equal utilization) that screens for instability from one model without retraining, and a multi-model attribution $Z$-test (F1: tests whether the attribution gap between features is significantly different from zero across $M$ models) that confirms instability with controlled type I error. The FSI and IS Plot below are exploratory diagnostics that visualize the instability landscape; the $Z$-tests provide formal hypothesis testing. The recommended workflow combines both: screen $\to$ confirm $\to$ resolve (via DASH) $\to$ audit (via FSI/IS Plot).
@@
Quantitative validation of FSI.
On synthetic data with known ground truth, FSI cleanly separates signal features (those in groups with $\beta_{g}>0$) from noise features ($\beta_{g}=0$) at every correlation level (Table 7). Signal features have consistently lower FSI (more stable attributions) than noise features, with the ratio declining from $0.31$ at $\rho=0$ to $0.25$ at $\rho=0.95$—indicating that FSI becomes a sharper discriminator as collinearity increases. The Spearman correlation between FSI and ground-truth importance magnitude is $\rho_{S}\approx-0.995$ ($p<0.001$) at every level, confirming that FSI ordering nearly perfectly recovers the true importance ordering without access to ground truth.
Table 7: FSI validation on synthetic data (50 repetitions per $\rho$). Mean FSI for signal features ($\beta_{g}>0$) vs. noise features ($\beta_{g}=0$). $\beta$ Spearman is the rank correlation between FSI and ground-truth importance magnitude across all 50 features.
$\rho$ Mean FSI (signal) Mean FSI (noise) Ratio $\beta$ Spearman
0.0 0.285 0.905 0.31 -0.995
0.5 0.337 1.253 0.27 -0.995
0.7 0.362 1.424 0.25 -0.995
0.9 0.420 1.636 0.26 -0.994
0.95 0.461 1.843 0.25 -0.991
@@ 6.5 Real-world validation
On the Breast Cancer dataset (30 features, 21 pairs with $|r|>0.9$, 50 repetitions), DASH nearly triples stability:
@@
Breast Cancer Single Best $376\pm.043$
Single Best (M=200M{=}200) $339\pm.037$ 0.037-0.037
Stochastic Retrain $862\pm.010$ +0.486+0.486
Random Selection $919\pm.004$ +0.543+0.543
Random Forest $922\pm.004$ +0.546+0.546
DASH (MaxMin) $\mathbf{.925}\pm.004$ +0.549+0.549
Superconductor Single Best $840\pm.014$ $9.22\pm 0.11$
Large Single Model $721\pm.008$ 0.119-0.119 $9.36\pm 0.12$
Stochastic Retrain $924\pm.008$ +0.084+0.084 $9.16\pm 0.09$
DASH (MaxMin) $964\pm.001$ +0.124+0.124 $9.17\pm 0.09$
Naive Top-N $\mathbf{.976}\pm.001$ +0.136+0.136 $\mathbf{9.15}\pm 0.10$
Calif. Housing Single Best $969\pm.003$ $0.459\pm 0.007$
Stochastic Retrain $977\pm.002$ +0.008+0.008 $0.450\pm 0.005$
DASH (MaxMin) $\mathbf{.978}\pm.004$ +0.009+0.009 $0.452\pm 0.004$
Random Forest $\mathbf{.998}\pm.001$ +0.029+0.029 $0.517\pm 0.004$
The Breast Cancer improvement is the largest across all experiments and the one real-world dataset where DASH clearly outperforms Stochastic Retrain ($0.925$ vs. $0.862$, +0.063+0.063)—the largest DASH-SR gap in any experiment. This contrasts with the synthetic linear sweep, where the two methods are statistically equivalent, and suggests that Breast Cancer’s overlapping chain correlations (the radius/perimeter/area triad) require the feature-level diversity that DASH’s forced low colsample_bytree and MaxMin selection provide. The training-budget-matched Single Best (M=200M{=}200) achieves only $0.339$, worse than the standard Single Best ($0.376$), because extreme collinearity makes model selection itself unstable—with 200 hyperparameter configurations, the “best” model varies wildly across repetitions. Random Forest achieves comparable stability ($0.922$) via internal feature bagging, but its near-zero ablation score ($0.005$ vs. DASH’s $0.143$) indicates that RF’s attributions are largely invariant to feature removal—stable through marginalization rather than through accurate feature-level sensitivity. DASH consensus produces stable top features aligned with known geometric redundancy—mean concave points (0.228), worst area (0.201), worst perimeter (0.201)—that are stable across repetitions. These rankings are consistent with the clinical literature: concavity features reflect nuclear contour irregularity, a hallmark of malignancy, while area and perimeter capture tumor size (Street et al., 1993).
@@
On the Superconductor dataset (81 features, 21,263 samples), DASH improves stability by +0.124+0.124 over Single Best ($0.964$ vs. $0.840$) and +0.243+0.243 over the Large Single Model, while achieving comparable predictive RMSE. However, Random Selection ($0.968$) and Naive Top-N ($0.976$) slightly exceed DASH ($p<0.001$ for both), suggesting that MaxMin diversity selection provides diminishing returns when the natural feature diversity is already high. With 81 features, most models in the population already produce distinct attribution profiles, reducing the value of explicit diversity maximization. The LSM result remains striking: despite matching DASH’s total tree count, its sequential training produces the poorest reproducibility (0.7210.721)—consistent with first-mover bias at scale.
@@
On California Housing (8 features, 20,640 samples), several feature pairs have $|r|>0.7$ (e.g., latitude/longitude, income/house value). DASH improves stability by +0.009+0.009 over Single Best, a difference that is not statistically significant ($p=0.063$). DASH and Stochastic Retrain are statistically equivalent (TOST confirmed), consistent with the milder degree of collinearity and the small feature space. Aggregation-based methods (Random Selection, Ensemble SHAP) achieve higher stability on this dataset, suggesting that simple model averaging suffices when collinearity is mild. Random Forest achieves the highest stability ($0.998$) but with substantially worse RMSE ($0.517$ vs. $0.452$).
@@ 6.6 Robustness and scope
Overlapping correlation structure.
The main synthetic DGP uses block-diagonal correlation. Real data rarely has such clean structure. To test robustness, we use a chain-correlation DGP where groups share features (A–B–C overlap). DASH achieves stability $0.976$ vs. Single Best $0.897$ (+0.079+0.079)—the largest stability advantage observed in any synthetic experiment—and also dominates on top-5 agreement (+0.156+0.156) and equity (0.067-0.067). This confirms that DASH’s MaxMin selection is particularly effective when the correlation structure is complex, as it selects models that explore different parts of the overlapping feature space.
@@
Under the nonlinear DGP (Table 9), all methods degrade (stability drops from $\sim$0.93 to $\sim$0.88), but the first-mover bias mechanism still operates: DASH outperforms Single Best at $\rho\geq 0.7$ (stability gap +0.083+0.083 at $\rho=0.95$).
Table 9: Nonlinear DGP: stability and equity across correlation levels (50 repetitions per $\rho$ level). SB = Single Best, LSM = Large Single Model, LSM-T = LSM (Tuned), SR = Stochastic Retrain. Five methods with distinct training structures are shown; Random Selection and Naive Top-$N$ (which performed similarly to DASH in the linear regime) are omitted for space and are available in the code repository. Bold indicates best per $\rho$.
@@
0.0 0.174 0.179 0.178 0.177 0.163
0.5 0.168 0.189 0.174 0.166 0.165
0.7 0.185 0.208 0.187 0.182 0.173
0.9 0.217 0.242 0.217 0.184 0.161
0.95 0.230 0.261 0.240 0.180 0.159
At $\rho=0$, DASH and Single Best perform nearly identically ($0.933$ vs. $0.931$). At $\rho=0.5$, Stochastic Retrain achieves the highest stability ($0.860$), with DASH ($0.857$) marginally ahead of Single Best ($0.850$). DASH’s advantage over Single Best grows consistently with correlation. This is a genuine scope boundary: nonlinear relationships are common in practice, and overall stability levels are lower than for the linear DGP ($\sim$0.86–0.93 vs. $\sim$0.93–0.98). When the DGP contains interactions and nonlinear terms, different models in the DASH ensemble may capture different interaction structures. Averaging SHAP values across models that have learned qualitatively different functional forms introduces noise rather than canceling arbitrary choices.
At $\rho\geq 0.7$, first-mover bias reasserts itself as the dominant source of instability, and DASH’s independence-based cancellation provides clear benefit (+0.083+0.083 at $\rho=0.95$). Practitioners should therefore check both the degree of collinearity and the presence of strong nonlinear interactions before applying DASH. The FSI diagnostic (Section 6.4) can help: if FSI values are low across all features, ensemble averaging may not be needed.
@@ 7.1 Reframing explanation instability
This work reframes the SHAP instability problem from “SHAP is noisy under multicollinearity” to a specific mechanistic hypothesis: sequential residual dependency in gradient boosting creates a first-mover bias that concentrates feature importance on arbitrary features within correlated groups. Three lines of evidence support this view: (1) the Large Single Model—which maximizes sequential dependency—produces the poorest reproducibility at every correlation level; (2) all methods that achieve model independence restore stability to the same level ($\approx 0.977$ at $\rho=0.9$), regardless of their other design choices; and (3) the effect scales with correlation severity, as the hypothesis predicts. The independence principle would be falsified by a setting in which averaging over independently trained models produces lower stability than a single model—a scenario we have not observed but cannot rule out under adversarial construction or extreme model heterogeneity (e.g., if independently trained models learn qualitatively different functions rather than merely different feature orderings).
@@
Random forests as an independence control.
Random Forest (RF) provides a natural test of the independence principle from the model-family side: its trees are trained independently by construction (via bootstrap sampling and random feature subsets). The companion impossibility analysis (Caraker et al., 2026b) predicts this theoretically: RF’s attribution ratio converges as $O(1/\sqrt{T})$, in contrast to gradient boosting’s divergent $1/(1-\rho^{2})$—independence between trees prevents the compounding that drives first-mover concentration. In the linear correlation sweep, RF achieves stability competitive with DASH ($0.978$ vs. $0.977$ at $\rho=0.9$), consistent with this prediction. However, RF’s top-$5$ ranking agreement is substantially lower ($0.375$ vs. $0.863$; $p<0.001$, $d=+0.488$) and its predictive RMSE is significantly worse ($0.957$ vs. $0.591$). This pattern—high stability but poor top-feature identification—indicates that RF produces stable but systematically different attributions. Independence ensures that arbitrary choices cancel upon averaging, yielding high stability; but if the base models are individually inaccurate (e.g., due to high prediction error), the stable consensus may converge to an incorrect ranking. DASH achieves both stability and accurate attributions because its base models are individually performant XGBoost learners filtered for predictive quality before aggregation.
@@
The DASH ensemble’s $K$ importance vectors also support a richer output representation: features that consistently swap rank across ensemble members can be treated as incomparable rather than arbitrarily ordered, yielding a partial order on feature importance that more honestly reflects genuine attribution ambiguity under collinearity. We pursue this direction in companion work (Caraker et al., 2026b), which formalizes the conditions under which partial orders are necessary rather than merely convenient.